It takes James and Bret 15 minutes to travel 5 miles upstream against the current. The return trip downstream with the current takes 10 minutes. The speed of the current remained constant.
I need help with finding their speed for upstream and downstream and then finding the boat and current speeds...

Answer by stanbon(75887)

It takes James and Bret 15 minutes to travel 5 miles upstream against the current. The return trip downstream with the current takes 10 minutes. The speed of the current remained constant.
-------
Upstream DATA:
distance = 5 miles ; time = 1/4 hr ; rate = d/t = 5/(1/4) = 20 mph
----
Downstream DATA:
distance = 5 miles ; time = 1/6 hr ; rate = d/t = 5/(1/6) = 30 mph
----
Equation:
b + c = 30 mph
b - c = 20 mph
----
Add and solve for "b":
2b = 50 mph
b = 25 mph (speed of the boat in still water)
---
Solve for "c":
b + c = 30
25 + c = 30
c = 5 mph (speed of the current)
====================================
Cheers,
Stan H.
==============